# Almost mathematics of pointed symmetric monoidal model categories by Smith ideal theory

### v4

@@ 5.3. Definition of operadic non-unital algebra objects of symmetric monoidal categories
Let $\mathcal{M}$ be a symmetric monoidal closed model category with a monoidal unit $V$. An homotopically augmented $\mathcal{O}$-algebra over $V$ $f:A\to B$ is an object of $\mathrm{Alg}_{\overrightarrow{\mathcal{O}^{\otimes}}}(\mathrm{Ar}^{\otimes}(\mathcal{M}))$ that the unit $U^{0}(V)=(\mathrm{id}_{V}:V\to V)\to(f:A\to B)$ has a homotopy section. Let $\mathrm{Alg}_{\mathcal{O}^{\otimes}}(\mathrm{Ar}^{\otimes}(\mathcal{M}))_{V//V}$ denotes the semi-model category of homotopically augmented $V$-algebras. In particular, $\mathrm{CAlg}(\mathcal{M})_{V//V}$ denotes the semi-model category of homotopically augmented $V$-algebras.
@@
Due to the following White and Yau’s theorem, we have a Quillen equivalence between semi-model categories of non-unital non-unital $\mathcal{O}$-algebras and homotopically augmented $\mathcal{O}$-algebras.
Theorem 5.8 ([WY24] Theorem 4.4.1).
@@
Definition 5.10 ([Whi22], Definition 3.1 and 3.4).
A symmetric monoidal model category $\mathcal{M}$ is said to be satisfying the commutative monoid axiom if, for all $n>0$, the $n$-fold symmetric pushout product functor $(-)^{\Box n}/\Sigma_{n}:\mathcal{M}\to\mathcal{M}$ preserves all trivial cofibrations. Further if the $n$-fold symmetric pushout product functor preserves all cofibrations and trivial cofibrations, $\mathcal{M}$ is said to be satisfying the strong commutative monoid axiom.
In particular, by [WY24, Corollary 5.1.2], we obtain the following:
@@
Let $\mathcal{M}$ be a cofibrantly generated stable model category satisfying the strong commutative monoid axiom, the monoid axiom, and in which cofibrant commutative $\mathrm{Ar}^{\Box}(\mathcal{M})$. Then the cokernel-and-kernel adjunction induces a Quillen equivalence
@@
If $\mathcal{M}$ is a locally presentable stable symmetric monoidal model category, the presentable $\infty$-category of augmented algebra is represented by the model category $\mathrm{CAlg}(\mathcal{M})_{V//V}$. Hence the left-hand-side $\mathrm{CAlg}^{\rm nu}(\mathcal{M})$ is equivalent to the $\infty$-category defined by Lurie [Lur17, p.949, Definition 5.4.4.9 and Proposition 5.4.4.10].

### v5

@@ 5.3. Definition of operadic non-unital algebra objects of symmetric monoidal categories
Let $\mathcal{M}$ be a symmetric monoidal closed model category with a monoidal unit $V$. An homotopically augmented $\mathcal{O}$-algebra over $V$ $f:A\to B$ is an object of $\mathrm{Alg}_{\overrightarrow{\mathcal{O}^{\otimes}}}(\mathrm{Ar}^{\otimes}(\mathcal{M}))$ such that the unit $U^{0}(V)=(\mathrm{id}_{V}:V\to V)\to(f:A\to B)$ has a homotopy section. Let $\mathrm{Alg}_{\mathcal{O}^{\otimes}}(\mathrm{Ar}^{\otimes}(\mathcal{M}))_{V//V}$ denote the semi-model category of homotopically augmented $V$-algebras. In particular, $\mathrm{CAlg}(\mathcal{M})_{V//V}$ denotes the semi-model category of homotopically augmented $V$-algebras.
@@
By the following theorem of White and Yau, we have a Quillen equivalence between semi-model categories of non-unital $\mathcal{O}$-algebras and homotopically augmented $\mathcal{O}$-algebras.
Theorem 5.8 ([18] Theorem 4.4.1).
@@
Definition 5.10 ([19], Definition 3.1 and 3.4).
A symmetric monoidal model category $\mathcal{M}$ is said to satisfy the commutative monoid axiom if, for all $n>0$, the $n$-fold symmetric pushout product functor $(-)^{\Box n}/\Sigma_{n}:\mathcal{M}\to\mathcal{M}$ preserves all trivial cofibrations. Further, if the $n$-fold symmetric pushout product functor preserves all cofibrations and trivial cofibrations, $\mathcal{M}$ is said to satisfy the strong commutative monoid axiom.
In particular, by [18, Corollary 5.1.2], we obtain the following:
@@
Let $\mathcal{M}$ be a cofibrantly generated stable model category satisfying the strong commutative monoid axiom and the monoid axiom, and assume that the category of commutative monoids in $\mathrm{Ar}^{\Box}(\mathcal{M})$ has the transferred model structure. Then the cokernel-and-kernel adjunction induces a Quillen equivalence
@@
If $\mathcal{M}$ is a locally presentable stable symmetric monoidal model category, the presentable $\infty$-category of augmented algebra is represented by $\mathrm{CAlg}(\mathcal{M})_{V//V}$. Hence the left-hand side $\mathrm{CAlg}^{\rm nu}(\mathcal{M})$ is equivalent to the $\infty$-category defined by Lurie [10, p.949, Definition 5.4.4.9 and Proposition 5.4.4.10].
Acknowledgements
I would like to thank anonymous referees for giving me innovative ideas and constructive comments to use Hovey’s Smith ideal theory and White and Yau’s operadic Smith ideal theory. I am also grateful to Isamu Iwanari for pointing out the derived-completion example.
Use of AI tools.
The author used Google’s Gemini-pro 3.1 as a conversational research aid for brainstorming and refining mathematical formulations, and OpenAI’s Prism (GPT-5.2) for editorial and expository assistance, including structural refinement. These tools were used to improve the manuscript’s clarity and to indicate where abbreviated arguments warranted fuller exposition. All results, proofs, and final formulations were reviewed, verified, and approved by the author, who bears sole responsibility for the content of this work.